# On multi-graded Proj schemes

### v4

@@ 5.4 Maximally relevant families
Recall that if $\mathcal{L}$ is a line bundle on a scheme $X$ and $s\in\Gamma(X,\mathcal{L})$ is a global section, we have an open subscheme $X_{s}\subset X$ defined by the nonvanishing of $s$, see [StP, Tag 01CY].
@@
Assume that (5.8) is satisfied. Let $f\in A$ be homogeneous and relevant, and let $\tilde{f}$ be the image of $f$ under the morphism of Fact 5.7(1). We have
@@
We can assume that $f$ is nonzero. Our assumptions imply that ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$ is a line bundle (see Corollary 5.11), so that ${\operatorname{Proj}}^{M}(A)_{\tilde{f}}$ is well defined. In view of our assumption, to prove the statement it suffices to prove that for any $S\in{\mathcal{F}}^{\mathrm{m}}_{A}$ we have
@@
$\Gamma(D_{\dagger}(f),\widetilde{\Gamma_{\bullet}({\mathcal{Q}})})\to\Gamma(D_{\dagger}(f),{\mathcal{Q}})$
@@
Under the additional assumption that $M$ is a free abelian group, one can give an alternative proof of Proposition 5.19 as follows. As explained in the comments following (3.4), the natural morphism ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Proj}}^{M}(A)$ is affine, hence quasi-compact, so that the scheme $A\smallsetminus V(A_{\dagger})$ is quasi-compact. Hence the open immersion $j:{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Spec}}(A)$ is quasi-compact and separated, so that the pushforward functor $j_{*}$ preserves quasi-coherent sheaves, see [StP, Tag 01LC]. Under the identifications considered in Remark 5.14 the functor $Q\mapsto\widetilde{Q}$ corresponds to $j^{*}$, while the functor $\Gamma_{\bullet}$ corresponds to $j_{*}$, and the statement of the proposition becomes the familiar fact that the adjunction morphism $j^{*}j_{*}\to{\operatorname{id}}$ is an isomorphism.
@@
$\mathsf{L}:{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}\to\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$
@@
$\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))\to{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})\xrightarrow{\Gamma_{\bullet}}{\mathsf{Mod}}^{M}(A)\to{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}$
@@
Let $\mathsf{A}$, $\mathsf{B}$ be abelian categories, and let $L:\mathsf{A}\to\mathsf{B}$ and $R:\mathsf{B}\to\mathsf{A}$ be functors. Assume that
@@
Then $L$ factors through an equivalence of categories $\overline{L}:\mathsf{A}/\ker(L)\xrightarrow{\sim}\mathsf{B}$, whose quasi-inverse is the composition of $R$ with the quotient functor $\mathsf{A}\to\mathsf{A}/\ker(L)$.
@@
We now consider the case when $A$ is noetherian. Recall that in this case the scheme ${\operatorname{Proj}}^{M}(A)$ is noetherian (in particular, quasi-compact), see §5.2. In this case we consider the category ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)$ of finitely generated $M$-graded $A$-module, its Serre subcategory ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)_{\mathrm{neg}}$ of objects which are negligible modules, and the category $\mathsf{Coh}({\operatorname{Proj}}^{M}(A))$ of coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. Recall (see Lemma 5.6) that in this case the functor $Q\mapsto\widetilde{Q}$ restricts to a functor ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)\to\mathsf{Coh}({\operatorname{Proj}}^{M}(A))$, which must factor through a functor
$\mathsf{L}_{\mathsf{Coh}}:{\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)/{\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)_{\mathrm{neg}}\to\mathsf{Coh}({\operatorname{Proj}}^{M}(A)).$
@@
${\mathsf{Mod}}^{M,H}(A)/{\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}\to\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$
@@
${\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)/{\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)_{\mathrm{neg}}\to\mathsf{Coh}^{H}({\operatorname{Proj}}^{M}(A))$
@@ 5.6 On a lemma by Arkhipov–Bezrukavnikov
In [AB09, Sublemma 1], the authors state a similar claim for bounded homotopy categories. [^6] Note that for a triangulated category $\mathsf{A}$ and full subcategories $\mathsf{B}$, $\mathsf{C}$, it is not true in general that the natural functor $\mathsf{B}/(\mathsf{B}\cap\mathsf{C})\to\mathsf{A}/\mathsf{C}$ is fully faithful; the bounded case therefore does not immediately follow from the unbounded case. The proof is very sketchy, and it is not clear to us if it really applies in the stated generality. Here we provide a complete argument for this claim (under additional assumptions), kindly explained to one of us by R. Bezrukavnikov.
@@
$K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)\to D^{\mathrm{b}}\mathsf{Coh}({\operatorname{Proj}}^{M}(A)).$
@@
    $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)\to D^{\mathrm{b}}\mathsf{Coh}^{H}({\operatorname{Proj}}^{M}(A))$

### v5

@@ 5.4 Maximally relevant families
Recall that if $\mathcal{L}$ is an invertible sheaf on a scheme $X$ and $s\in\Gamma(X,\mathcal{L})$ is a global section, we have an open subscheme $X_{s}\subset X$ defined by the nonvanishing of $s$, see [StP, Tag 01CY].
@@
Assume that (5.8) is satisfied. Let $f\in A$ be homogeneous and relevant, and let $\tilde{f}$ be the image of $f$ in $\Gamma({\operatorname{Proj}}^{M}(A),{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$ (cf. Fact 5.7(1)). We have
@@
We can assume that $f$ is nonzero. Our assumptions imply that ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$ is an invertible sheaf (see Corollary 5.11), so that ${\operatorname{Proj}}^{M}(A)_{\tilde{f}}$ is well defined. In view of our assumption, to prove the statement it suffices to prove that for any $S\in{\mathcal{F}}^{\mathrm{m}}_{A}$ we have
@@
$\Gamma(D_{\dagger}(f),\widetilde{\Gamma_{\bullet}({\mathcal{Q}})})\longrightarrow\Gamma(D_{\dagger}(f),{\mathcal{Q}})$
@@
Under the additional assumption that $M$ is a free abelian group, one can give an alternative proof of Proposition 5.19 as follows. As explained in the comments following (3.4), the natural morphism ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Proj}}^{M}(A)$ is affine, hence quasi-compact, so that the scheme $A\smallsetminus V(A_{\dagger})$ is quasi-compact. Hence the open immersion $j:{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Spec}}(A)$ is quasi-compact and separated, so that the pushforward functor $j_{*}$ preserves quasi-coherent sheaves, see [StP, Tag 01LC]. Under the identifications considered in Remark 5.14, the functor $Q\mapsto\widetilde{Q}$ corresponds to $j^{*}$, while the functor $\Gamma_{\bullet}$ corresponds to $j_{*}$, and the statement of the proposition becomes the familiar fact that the adjunction morphism $j^{*}j_{*}\to{\operatorname{id}}$ is an isomorphism.
@@
$\mathsf{L}:{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}\longrightarrow\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$
@@
$\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))\longrightarrow{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})\xrightarrow{\Gamma_{\bullet}}{\mathsf{Mod}}^{M}(A)\longrightarrow{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}$
@@
Let $\mathsf{A}$, $\mathsf{B}$ be abelian categories, and let $L:\mathsf{A}\longrightarrow\mathsf{B}$ and $R:\mathsf{B}\longrightarrow\mathsf{A}$ be functors. Assume that
@@
Then $L$ factors through an equivalence of categories $\overline{L}:\mathsf{A}/\ker(L)\xrightarrow{\sim}\mathsf{B}$, whose quasi-inverse is the composition of $R$ with the quotient functor $\mathsf{A}\longrightarrow\mathsf{A}/\ker(L)$.
@@
We now consider the case when $A$ is noetherian. Recall that in this case the scheme ${\operatorname{Proj}}^{M}(A)$ is noetherian (in particular, quasi-compact), see §5.2. We consider the category ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)$ of finitely generated $M$-graded $A$-module, its Serre subcategory ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)_{\mathrm{neg}}$ of objects which are negligible modules, and the category $\mathsf{Coh}({\operatorname{Proj}}^{M}(A))$ of coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. Recall (see Lemma 5.6) that in this case the functor $Q\mapsto\widetilde{Q}$ restricts to a functor ${\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)\to\mathsf{Coh}({\operatorname{Proj}}^{M}(A))$, which must factor through a functor
$\mathsf{L}_{\mathsf{Coh}}:{\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)/{\mathsf{Mod}}^{M}_{\mathrm{fg}}(A)_{\mathrm{neg}}\longrightarrow\mathsf{Coh}({\operatorname{Proj}}^{M}(A)).$
@@
${\mathsf{Mod}}^{M,H}(A)/{\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}\longrightarrow\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$
@@
${\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)/{\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)_{\mathrm{neg}}\longrightarrow\mathsf{Coh}^{H}({\operatorname{Proj}}^{M}(A))$
@@ 5.6 On a lemma by Arkhipov–Bezrukavnikov
In [AB09, Sublemma 1], the authors state a similar claim for bounded homotopy categories. [^7] Note that for a triangulated category $\mathsf{A}$ and full subcategories $\mathsf{B}$, $\mathsf{C}$, it is not true in general that the natural functor $\mathsf{B}/(\mathsf{B}\cap\mathsf{C})\to\mathsf{A}/\mathsf{C}$ is fully faithful; the bounded case therefore does not immediately follow from the unbounded case. The proof is very sketchy, and it is not clear to us if it really applies in the stated generality. Here we provide a complete argument for this claim (under additional assumptions), kindly explained to us by R. Bezrukavnikov.
@@
$K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)\longrightarrow D^{\mathrm{b}}\mathsf{Coh}({\operatorname{Proj}}^{M}(A)).$
@@
    $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)\longrightarrow D^{\mathrm{b}}\mathsf{Coh}^{H}({\operatorname{Proj}}^{M}(A))$